# Text-Speech Language Models with Improved Cross-Modal Transfer by Aligning Abstraction Levels

### v1

@@ Abstract
Text-Speech Language Models (TSLMs)—language models trained to jointly process and generate text and speech—aim to enable cross-modal knowledge transfer to overcome the scaling limitations of unimodal speech LMs. The predominant approach to TSLM training expands the vocabulary of a pre-trained text LM by appending new embeddings and linear projections for speech, followed by fine-tuning on speech data. We hypothesize that this method limits cross-modal transfer by neglecting feature compositionality, preventing text-learned functions from being fully leveraged at appropriate abstraction levels. To address this, we propose augmenting vocabulary expansion with modules that better align abstraction levels across layers. Our models, SmolTolk, rival or surpass state-of-the-art TSLMs trained with orders of magnitude more compute. Representation analyses and improved multimodal performance suggest our method enhances cross-modal transfer.
Multimodal Language Models, Text-Speech Language Models, Speech Language Models, Language Models, Modality Alignment
@@ 1 Introduction
Generative auto-regressive language modeling has proven to be a powerful tool well beyond text. Since the paradigm-shifting arrival of Large Language Models (LLMs), a highly prolific line of research has focused on applying the same type of model to discrete representations (tokens) of other modalities, such as images (Chen et al., 2020), video (Yu et al., 2024), and audio (Lakhotia et al., 2021; Borsos et al., 2023). This unified approach has enabled the development of generative multimodal LMs, capable of jointly modeling different modalities (Aghajanyan et al., 2022; Chameleon Team, 2024; Rubenstein et al., 2023), paving the way for AIs that learn patterns across data of diverse nature and transfer knowledge between them (Aghajanyan et al., 2023). Given the remarkable capacities of text LLMs, a key goal is to enable cross-modal knowledge transfer between them and other modalities.
Because of its high mutual information with text, speech is well-suited for cross-modal transfer from text LMs. At the same time, there is a strong incentive to pursue this approach: unimodal speech LMs scale poorly, requiring orders of magnitude more compute to reach comparable linguistic competence (Cuervo & Marxer, 2024). This inefficiency suggests that simply increasing model and dataset sizes is unlikely to be viable, making cross-modal transfer not just promising but perhaps necessary. Accordingly, several works have explored joint text-speech modeling through Text-Speech Language Models (TSLMs) (Nguyen et al., 2024; Zhang et al., 2023; Défossez et al., 2024; Zeng et al., 2024), most following a common approach: expanding a pre-trained text LM’s vocabulary to accommodate speech by appending new embeddings and linear projections at the input and output respectively, then training on mixed speech and text data. While these methods demonstrate positive transfer, they remain heavily dependent on large amounts of paired text-speech data and still show a considerable performance gap between text and speech.
In this paper, we hypothesize that a key limitation of previous methods is their failure to account for the hierarchy of feature compositionality in text and speech. We note that: (1) Speech tokens, which encode acoustic and phonetic information, are at a lower level of abstraction than the word-like [^1] Assuming the widely used sub-word tokenization. tokens expected at the input and output of text LMs, as spoken words consist of multiple speech tokens. (2) LMs extract increasingly abstract features across layers, with the last few layers refining higher-level features into lower-level ones predictive of next tokens (Valeriani et al., 2023; Cheng et al., 2025). Thus, we argue that:
  1. 1.
    Since cross-modal transfer aims to enable functions learned during text pre-training to process speech’s linguistic information, speech tokens should first be composed into the word-like features these functions expect as input.
  2. 2.
    The output representations of text LMs are optimized to predict next words. Next-word information is useful for predicting next-speech tokens, and cross-modal transfer should leverage it. However, since next-speech tokens are at a lower level of abstraction than next words, the text LM’s output representations should be refined to predict next-speech tokens.
  3. 3.
    In speech language modeling, next-word information is most useful near word boundaries (or plausible boundaries), which are sparse relative to in-word tokens. Most of the time (within words), the information in the last layers for next-speech token prediction pertains to the current word and speech token—features that should be abstracted in the earlier layers. Therefore, we hypothesize that regular vocabulary expansion, which predicts speech tokens solely from text LM output representations, encourages the model to retain low-level information in later layers, possibly preventing it from abstracting higher-level features.
We posit that the linear functions used in vocabulary expansion methods are ill-suited for addressing points 1 and 2 and lack mechanisms to mitigate point 3. In this work, we test our hypotheses and aim to address these potential limitations. Our main contributions are:
  • We augment vocabulary expansion to better align abstraction levels across the model’s layers. To address points 1 and 2, we introduce additional speech-specific layers at the model’s input and output. To address point 3, we propose a learnable dynamic layer pooling mechanism that allows the speech output layers to adaptively switch between attending to low-level and next-word predictive representations.
@@
  • We apply our method to the SmolLM family of language models (Ben Allal et al., 2024), which range in size from 135 million to 1.7 billion parameters. The resulting models, named SmolTolk, consistently outperform baselines trained with regular vocabulary expansion. Our largest model, SmolTolk-2B, rivals or surpasses orders-of-magnitude larger TSLMs.
  • We analyze the contributions of different architectural components through ablations and representation analyses. Our results suggest that each proposed element improves overall performance and contributes to facilitate cross-modal transfer.
@@ 2 Text-Speech Language Models
Text-Speech Language Models (TSLMs) model the joint probability of text and speech token sequences as
@@
Text tokens are typically obtained via sub-word tokenization (Sennrich et al., 2016), while speech tokens are derived through quantization of self-supervised representations. Linguistic tokens [^2] Often referred to in the literature as ”semantic tokens”, though we argue this is a misnomer as they primarily encode phonetic information (Choi et al., 2024). are obtained from masked language modeling models like HuBERT (Hsu et al., 2021) and capture phonetic content. Acoustic tokens, extracted from autoencoder models, preserve speech signal details, including paralinguistic and acoustic variability (Zeghidour et al., 2021).
Training. TSLMs are trained via vocabulary expansion and speech fine-tuning of text LMs. Vocabulary expansion extends the embedding function and output projections over $\mathcal{V}_{t}$ to include $\mathcal{V}_{s}$, while the rest of the LM remains unchanged. Fine-tuning methods vary in data mixture: Rubenstein et al. (2023) train on mixed speech-text tasks (TTS, ASR, speech-to-speech translation), while Chou et al. (2023) use word-level alignments to switch modalities within a sequence. This interleaved text-speech strategy was shown to be crucial for cross-modal transfer, later validated and scaled up by Nguyen et al. (2024) and Zeng et al. (2024), achieving state-of-the-art speech LM performance.
Figure 1: General diagram of our proposed architecture. Input tokens are processed through modality-specific embedding tables. Speech representations (blue) pass through speech-specific layers, bridging the gap between speech and text inputs, before merging with text embeddings (green) in the text LM backbone. A weighted average of the backbone’s representations, computed using fixed and dynamic learned weights yields a representations better suited for speech prediction, which is processed through output speech-specific layers to predict the next speech tokens. In parallel, text tokens are predicted from the final text LM representation.
3 Our Method
Our method is illustrated in Figure 1. Below, we describe our model’s architecture and its training process.
@@ 3.1 Model
We apply the embedding function $E$ to the input sequence composed of text and speech tokens $\mathbf{w}$, yielding a sequence of text and speech embeddings $(E(w_{1}),\dots,E(w_{n}))=(\mathbf{z}_{1},\dots,\mathbf{z}_{n}$).
@@
Input speech adapter. Contiguous chunks of speech embeddings are processed through an input adapter $A_{in}:\mathbb{R}^{d}\to\mathbb{R}^{d}$, a stack of decoder transformer layers. The input adapter is meant to compose the speech token embeddings into higher-level representations, which we believe should facilitate cross-modal transfer by matching the abstraction level expected by the text LM input. For a contiguous chunk of speech embeddings $(\mathbf{z}_{i},\dots,\mathbf{z}_{i+k})$, the input adapter outputs a sequence $(\mathbf{z}^{\prime}_{i},\dots,\mathbf{z}^{\prime}_{i+k})$.
@@
The output of the input adapter and the text embeddings are fed into the text LM transformer at their respective positions in the input. For instance, given the input sequence $(w_{1},w_{2},w_{3},w_{4},w_{5})$, where only the third element is a text token, the sequence passed to the transformer layers after the adapter would be $(\mathbf{z}_{1},\mathbf{z}_{2},E(w_{3}),\mathbf{z}_{4},\mathbf{z}_{5})$. For each transformer layer $l$, we obtain a sequence of contextual representations $(\mathbf{c}_{1}^{(l)},\dots,\mathbf{c}_{n}^{(l)})$. As in regular text LMs and TSLMs, the text output logits are computed by applying $U$ to the contextual representations $\mathbf{c}_{i}^{(L)}$ at the last layer $L$. The speech output logits are computed as described next.
@@
Dynamic layer pooling and speech input residual. We believe speech language modeling requires the model to switch between two modes of operation depending on whether a word is being generated or a new word is about to start. Within a word, the model should use low-level representations encoding the spoken word and the current speech token within that word, as these features fully determine the next speech token to be emitted. When generating a new word, the model should instead rely on representations predictive of upcoming words, such as those in the later layers of the text LM. To enable this behavior, we use a learnable mechanism that attends to different layers’ representations in an input-dependent manner. A linear layer selector $S:\mathbb{R}^{d}\times\mathbb{R}^{L}$ maps a contextual representation $\mathbf{c}_{i}^{\prime}$ to a vector of weights $\boldsymbol{\omega}_{i}=(\omega_{i}^{(1)},\dots,\omega_{i}^{(L)})$. These weights are softmax-normalized and used to compute an input-dependent weighted average yielding a contextual multi-level representation $\bar{\mathbf{c}}_{i}$:
$\boldsymbol{\omega}_{i}$ $=\text{Softmax}(S(\mathbf{c}_{i}^{\prime}))$ (3)
$\bar{\mathbf{c}}_{i}$ $=\sum_{l=1}^{L}\omega_{i}^{(l)}\mathbf{c}_{i}^{(l)}$
@@
A key question is which representation to use as the layer selector’s input $\mathbf{c}_{i}^{\prime}$. We found that last layer’s representations $\mathbf{c}_{i}^{(L)}$ were not well suited as they often resulted in the selector collapsing to select a single layer. Rather than searching for the best layer—which would likely vary across different architectures—we use a weighted average of the contextual representations with learned input-independent weights:
@@
To provide information about the current speech token, we add a residual connection from the speech input embeddings to the multi-level contextual representation:
@@
$\bar{\mathbf{c}}_{i}^{\prime}$ contains both the information selected by layer pooling from the text LM layers and the current speech token.
Output speech adapter. The output adapter $A_{out}:\mathbb{R}^{d}\to\mathbb{R}^{d}$, a stack of decoder transformer layers, takes $\bar{\mathbf{c}}_{i}^{\prime}$ as input and refines it into a representation predictive of upcoming speech tokens, upon which the speech output logits are computed by applying $U$.
@@ 3.2 Training
We train our model to optimize Equation 2 on sequences from a data mixture similar to that proposed by Chou et al. (2023), including unimodal speech and text samples, as well as interleaved text-speech samples. Unlike previous works, we do not include ASR or TTS samples.
@@
Preventing layer selector collapse. In larger models, we observed that the layer selector $S$ sometimes collapses early in training, attending to a single layer. To mitigate this, we add an entropy maximization term to the loss to encourage diversity in its output:
@@ 4.1 Models and training
We use SmolLM models (Ben Allal et al., 2024) as text LM backbones, available in three sizes: 135 million, 360 million, and 1.7 billion parameters. We refer to the models resulting from applying our method to the SmolLM backbones as SmolTolk-150M, SmolTolk-400M and SmolTolk-2B. Following Nguyen et al. (2024), and to improve the modeling of long-range dependencies, we increase the RoPE base frequency from 10,000 to 100,000.
The input and output adapters consist of transformer layers matching the text backbone architecture. In initial experiments, we observed that using more than two adapter layers led to only marginal or no improvements; thus, we use two layers for all our models. This aligns with findings from Turetzky & Adi (2024) on adapting speech representations to frozen text LMs. Across models, adding the adapters introduces a parameter overhead between 11% and 15%.
For speech tokenization, we follow Hassid et al. (2023), using the same tokenizer, which quantizes HuBERT representations extracted at 25 Hz into a 500-token vocabulary. As is common in speech LM training on linguistic tokens, we collapse adjacent token repetitions.
All LMs are optimized with AdamW (Loshchilov & Hutter, 2019) and a weight decay of 0.1. Models use a constant learning rate of 3e-4 for text and backbone parameters in the 135M, 360M, and 1.7B baselines, and 1e-4 for SmolTolk-2B. SmolTolk models apply a trapezoidal schedule for speech-specific parameters in stage 1 (1B tokens, 1000 steps), with a 100-step linear warmup and a maximum learning rate 10× that of text parameters, decaying over the final 20% of stage 1. For SmolTolk-400M and SmolTolk-2B, we set $\beta=0.01$ in Equation 6.
We use a batch size of 1 million tokens with the full 2048-token context. Each batch contains equal proportions of speech, text, and interleaved text-speech. All models are trained for 16 billion tokens, except the largest models, which are trained for up to 32 billion.

### v2

# Late Fusion and Multi-Level Fission Amplify Cross-Modal Transfer in Text-Speech LMs

@@ Abstract
Text-Speech Language Models (TSLMs)—language models trained to jointly process and generate text and speech—are commonly trained through an early modality fusion/fission approach, in which both modalities are fed and predicted from a shared backbone via linear layers. We hypothesize that this approach limits cross-modal transfer by neglecting feature compositionality—specifically, the finer-grained nature of speech representations compared to text—preventing the emergence of a shared feature hierarchy within model layers. In this paper, we argue that this limitation can be addressed through late fusion and fission, with a fission process that accesses both high- and low-level features for speech generation. Our models implementing these principles, SmolTolk, rival or surpass state-of-the-art TSLMs trained with orders of magnitude more compute, and achieve significantly improved cross-modal performance relative to early fusion/fission baselines. Representation analyses further suggest that our method enhances the model’s ability to abstract higher-level, more semantic features from speech, and leads to increasingly shared representation spaces across layers.
@@ 1 Introduction
Generative auto-regressive language modeling has proven to be a powerful method well beyond the textual domain. Since the paradigm-shifting emergence of Large Language Models (LLMs), a highly prolific line of research has focused on extending this modeling approach to discrete representations (tokens) of other modalities, such as images (Chen et al., 2020), video (Yu et al., 2024a), and audio (Lakhotia et al., 2021; Borsos et al., 2023). This unified framework has enabled the development of multimodal generative LMs capable of jointly modeling multiple modalities (Aghajanyan et al., 2022; Chameleon Team, 2024; Rubenstein et al., 2023), paving the way for AI systems that learn patterns across heterogeneous data and transfer knowledge between modalities (Aghajanyan et al., 2023).
Given the remarkable capabilities of text LLMs and the relative abundance of high-quality textual datasets compared to other modalities, enabling cross-modal knowledge transfer from text LLMs to lower-resource modalities has become a central goal. A common approach is to fine-tune pretrained text LLMs on multimodal data. However, this raises a fundamental question: How should a model originally designed to process and generate text adapt to handle non-textual modalities so as to promote cross-modal knowledge transfer? On the input side, this challenge is known as modality fusion—the problem of integrating different modalities into a shared representation space. On the output side, a related but less studied challenge arises: multimodal fission [^1] Perhaps a non-standard term, but with precedents in the literature (Landragin, 2007; Marinov et al., 2023)., which refers to the process of transforming a shared multimodal representation into modality-specific outputs.
Methods for modality fusion are typically categorized as either late or early fusion. In late fusion, non-textual modalities are first encoded by specialized modality-specific encoders—such as Whisper for speech (Yu et al., 2024a) or CLIP for images (Liu et al., 2023)—and optionally aligned using modules such as the Q-Former (Li et al., 2023), before their representations are passed to the text LLM. This approach was common in early multimodal LLMs designed for perception tasks with text-only generation. However, these architectures often rely on continuous representations from bidirectional encoders, which are not suitable for causal language modeling. More recent models for multimodal generation—sometimes referred to as omnimodal—have largely adopted early fusion instead, where non-textual modalities are fed directly in their native discrete representations (Chen et al., 2020), or as vector-quantized embeddings from self-supervised models not explicitly aligned to the text input space (Chameleon Team, 2024; Nguyen et al., 2024). These models typically adopt early fission, predicting modality-specific outputs linearly from the shared transformer backbone.
Early fusion/fission models benefit from a simpler, unified, and more end-to-end architecture. However, this comes at the cost of burdening the language model backbone with the task of internal modality alignment. In this paper, we hypothesize that early fusion and fission architectures are fundamentally ill-suited for effective cross-modal knowledge transfer. Our argument rests on three key observations:
  1. (i)
    Cross-modal transfer from text LLMs aims to leverage functions learned during text pretraining to process shared, “text-like” structure in non-textual modalities—such as linguistic content in speech or semantic scene descriptions in images.
  2. (ii)
    Textual features (e.g., words or subwords) are inherently higher-level abstractions than features in other modalities. A single text token can denote a complex image object composed of many pixels or patches, or a spoken word be composed of many discrete speech tokens, which are usually phonetic or sub-phonetic in nature.
  3. (iii)
    Language models build increasingly compositional features across layers, with deeper layers refining these abstractions into lower-level features optimized for next-token prediction (Valeriani et al., 2023; Cheng et al., 2025).
From these premises, we argue:
  1. (a)
    To fulfill (i), non-textual inputs must reach the level of abstraction expected by pretrained text functions—which, per (iii), increases with depth. Given (ii), this requires composing low-level modality-specific features into higher-level representations. However, in early fusion, all modalities pass through the same number of layers despite starting at different abstraction levels, leaving non-textual features under-composed. For instance, early layers that combine words into part-of-speech units in text would first need to build syllabic or word-level structures from speech tokens. This mismatch suggests the need for an input adapter that performs additional composition prior to the text backbone—i.e., late fusion.
  2. (b)
    The text LM output layer is tuned to predict next word-like tokens. These predictions contain valuable information for prediction in other modalities—e.g., the next speech tokens depend on the next word, or the next image patches may correlate with it. However, as per (ii), non-textual targets are typically of lower abstraction. To predict them from backbone outputs, a form of decomposition is needed—mirroring how word-level predictions are refined from more abstract representations (iii). This motivates an output adapter that transforms LM outputs into low-level non-textual targets—we call such process late fission.
  3. (c)
    Because of (ii), text-like features underlying non-textual modalities vary less frequently across time or space than non-textual features. As a result, next-word predictive features at the output of the backbone—optimized to model text dynamics—are likely most informative at sparse intervals (e.g., near word boundaries in speech). For much of the input (e.g., within a spoken word), non-textual next-token prediction relies instead on lower-level features abstracted earlier in the network. For example, given the speech utterance ‘‘we submitted the paper to the ’’, predicting upcoming speech tokens initially benefits from high-level features (e.g., anticipating ‘‘conference’’). However, while generating ‘‘conference’’, accurate prediction depends on fine-grained context—e.g., knowing the current phone ‘f‘ is necessary to predict the next phone ‘e‘. This suggests the need for a fission mechanism that accesses both low- and high-level features, possibly in an input-dependent manner. We refer to such a mechanism as multi-level fission.
In this work, we test these hypotheses in the context of text-speech language models. We focus on speech due to its high mutual information with text—making it especially suitable for cross-modal transfer from text-based LMs—and the prevalence of early fusion/fission approaches in the field (Hassid et al., 2023; Nguyen et al., 2024; Défossez et al., 2024; Zeng et al., 2024). At the same time, there is strong practical motivation: unimodal speech LMs scale poorly, requiring orders of magnitude more compute to reach comparable linguistic competence (Cuervo and Marxer, 2024). This inefficiency suggests that simply scaling models and data may not be viable, making cross-modal transfer not only promising but perhaps necessary for achieving high-quality speech generation.
Our main contributions are:
  • We propose an architecture featuring late fusion and late fission via simple causal adapters, and multi-level fission through an attention-like mechanism across layers, combined with a residual connection from early to late layers.
@@
  • We apply this architecture to the SmolLM family of text LMs (Allal et al., 2025), ranging from 135M to 1.7B parameters. The resulting models, SmolTolk, consistently outperform early fusion/fission baselines. Our largest variant, SmolTolk-2B, rivals or surpasses state-of-the-art models trained with orders of magnitude more compute.
  • We conduct ablations and representational analyses to isolate the impact of each architectural component. Results confirm that all proposed elements contribute to improved cross-modal transfer and performance, and reveal representation dynamics that support our hypotheses.
@@ 2 Text-Speech Language Models
TSLMs model the joint probability of text and speech token sequences as
@@
Training. TSLMs are typically trained via vocabulary expansion and speech fine-tuning of text LMs. Vocabulary expansion extends the embedding function and output projections over $\mathcal{V}_{t}$ to include $\mathcal{V}_{s}$, while the rest of the LM remains unchanged. Fine-tuning methods vary in data mixture: Rubenstein et al. (2023) train on mixed speech-text tasks (TTS, ASR, speech-to-speech translation), while Chou et al. (2023) use word-level alignments to switch modalities within a sequence. This interleaved text-speech strategy was shown to be crucial for cross-modal transfer, later validated and scaled up by Nguyen et al. (2024) and Zeng et al. (2024), achieving state-of-the-art speech LM performance.
Figure 1: Proposed architecture. The model processes interleaved text-speech sequences. The [swt] token denotes a modality switch. Late fusion: Speech inputs (blue) are processed by speech-specific layers before merging with text embeddings (green) in the text LM backbone. Multi-level fission: an input speech residual and an average across layers’ representations with input-dependent weights produce multi-layer representations. Late fission: These are passed through output speech-specific layers to predict speech tokens. Text tokens are predicted from the final backbone layer.
3 Proposed method
Our proposed method is illustrated in Figure 1, and described in detail below.
@@ 3.1 Model
The embedding function $E$ is applied to the input sequence composed of text and speech tokens $\mathbf{w}$, yielding a sequence of text and speech embeddings $(E(w_{1}),\dots,E(w_{n}))=(\mathbf{z}_{1},\dots,\mathbf{z}_{n}$).
@@
Late fusion. Contiguous chunks of speech embeddings are processed through an input adapter $A_{\text{in}}:\mathbb{R}^{d}\to\mathbb{R}^{d}$, designed to compose speech token embeddings into higher-level representations suitable as input to the text backbone. Unlike most late fusion architectures, which typically rely on Whisper-like pre-trained bidirectional encoders followed by downsampling and non-linear projections (Yu et al., 2024b; Chu et al., 2023), we employ a simple stack of decoder-only transformer layers applied directly to the speech tokens, maintaining a unified input/output representation. For a contiguous chunk of speech embeddings $(\mathbf{z}_{i},\dots,\mathbf{z}_{i+k})$, the input adapter outputs a sequence $(\mathbf{z}^{\prime}_{i},\dots,\mathbf{z}^{\prime}_{i+k})$.
@@
The output of the input adapter and the text embeddings are fed into the text LM transformer at their respective positions in the input. For instance, given the input sequence $(w_{1},w_{2},w_{3},w_{4},w_{5})$, where only the third element is a text token, the sequence passed to the transformer layers after the adapter would be $(\mathbf{z}_{1}^{\prime},\mathbf{z}_{2}^{\prime},E(w_{3}),\mathbf{z}_{4}^{\prime},\mathbf{z}_{5}^{\prime})$. For each transformer layer $l$, a sequence of contextual representations $(\mathbf{c}_{1}^{(l)},\dots,\mathbf{c}_{n}^{(l)})$ is obtained. As in standard text LMs and TSLMs, the text output logits are computed by applying $U$ to the contextual representations $\mathbf{c}_{i}^{(L)}$ at the last layer $L$. The speech output logits are computed as described next.
@@
Multi-level fission. As stated in Section 1, hypothesis (c), we posit that speech language modeling requires the model to alternate between two modes of operation: one for generating speech within a word, and another for initiating a new word. Within a word, the model should rely on low-level representations encoding the current word and its internal speech token, as these determine the next speech token. When starting a new word, the model should instead draw on higher-level features predictive of upcoming words, such as those found in the later layers of the text LM. To support this behavior, we introduce a learnable mechanism that attends to representations from different layers in an input-dependent manner. A linear layer selector $S:\mathbb{R}^{d}\to\mathbb{R}^{L}$ maps a contextual representation $\mathbf{c}_{i}^{\prime}$ to a vector of weights $\boldsymbol{\omega}_{i}=(\omega_{i}^{(1)},\dots,\omega_{i}^{(L)})$, which are softmax-normalized and used to compute a weighted average over layer representations, yielding the contextual multi-level representation $\bar{\mathbf{c}}_{i}$:
$\boldsymbol{\omega}_{i}=\text{Softmax}(S(\mathbf{c}_{i}^{\prime}))\hskip 19.91684pt,\hskip 19.91684pt\bar{\mathbf{c}}_{i}=\sum_{l=1}^{L}\omega_{i}^{(l)}\mathbf{c}_{i}^{(l)}$ (3)
@@
A key question is which representation to use as $\mathbf{c}_{i}^{\prime}$. We found that the last layer representations $\mathbf{c}_{i}^{(L)}$ are not well suited as they often resulted in the selector collapsing to select a single layer. Rather than searching for the best layer—which would likely vary across different architectures—we use a weighted average of the contextual representations with learned input-independent weights:
@@
To provide information about the current speech token, a residual connection from the speech input embeddings is added to the multi-level contextual representation:
@@
Here, $\bar{\mathbf{c}}_{i}^{\prime}$ contains both the layer-pooled features and the current speech token.
Late fission. The output adapter $A_{out}:\mathbb{R}^{d}\to\mathbb{R}^{d}$, a stack of decoder transformer layers, takes $\bar{\mathbf{c}}_{i}^{\prime}$ as input and refines it into a representation predictive of upcoming speech tokens, upon which the speech output logits are computed by applying $U$.
@@ 3.2 Training
We train our model to optimize Eq. 2 on a data mixture similar to Chou et al. (2023), including unimodal speech, text, and interleaved text-speech samples. Unlike prior work, we exclude ASR and TTS samples (except in the experiment in Appendix D.4).
@@
Preventing layer selector collapse. In larger models, we observed that the layer selector $S$ sometimes collapses early in training, attending to a single layer. To mitigate this, an entropy maximization term is added to the loss to encourage diversity in the model layer selection:
@@ 4.1 Models and training
We use SmolLM models (Allal et al., 2025) as text LM backbones, available in three sizes: 135 million, 360 million, and 1.7 billion parameters. We refer to the models resulting from applying our method to the SmolLM backbones as SmolTolk-150M, SmolTolk-400M and SmolTolk-2B.
The input and output adapters consist of transformer layers matching the text backbone architecture. In initial experiments, we observed that using more than two adapter layers led to only marginal or no improvements; thus, we use two layers for all our models. This aligns with findings from Turetzky and Adi (2024) on adapting speech representations to frozen text LMs. Across models, adding the adapters introduces a parameter overhead between 11% and 15%.
For speech tokenization, we follow Hassid et al. (2023), using the same tokenizer, which quantizes HuBERT representations extracted at 25 Hz into a 500-token vocabulary. As is common in speech LM training, adjacent token repetitions are collapsed to a single token.
All LMs are optimized with AdamW (Loshchilov and Hutter, 2019) and use a weight decay of 0.1. Models use a constant learning rate of 3e-4 for text and backbone parameters in the 135M, 360M, and 1.7B baselines, and 1e-4 for SmolTolk-2B. SmolTolk models apply a trapezoidal schedule for speech-specific parameters in stage 1 (1B tokens, 1000 steps), with a 100-step linear warmup and a maximum learning rate 10× that of text parameters, decaying over the final 20% of stage 1. For SmolTolk-400M and SmolTolk-2B, we set $\beta=0.01$ in Equation 6. We use a batch size of 1 million tokens with the full 2048-token context. Each batch contains equal proportions of speech, text, and interleaved text-speech. All models are trained for 16 billion tokens, except the largest models, which are trained for up to 32 billion.